Find f(a) if f(t) = t2 – 2t – 2
Answer by luke94(89)

hi there f(t) = t2 – 2t – 2 is a function or equation where t is just the variable being used.
your question is find f(a) meaning replace t with a.
f(a) = a^2 - 2a - 2
another example would be say if you were to find f(3):
f(3) = 3^2 - 2(3) - 2
= 9 - 6 - 2 = 1
hope this helps